# Exponential actions defined by vector configurations, Gale duality, and moment-angle manifolds

### v1

@@ 13. LVM and LVMB manifolds
We start with a spanning vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$ for which there exists ${\textbf{{v}}}_{1}\in V$ such that $\langle\gamma_{i},{\textbf{{v}}}_{1}\rangle=1$ for $i=1,\ldots,m$. Equivalently, the Gale dual configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ satisfies the relation $\sum_{k=1}^{m}{\textbf{{a}}}_{i}=\bf 0$. In particular, the conditions of Propostion 9.1 are met. Let $R=\mathbb{R}\langle{\textbf{{v}}}_{1}\rangle$. Thinking in terms of relations between the vectors of $\mathrm{A}$, the subspace $R$ is generated by a single relation with all coefficients $1$, so it is rational. The lattice (11.1) is the $\mathbb{Z}$-span of ${\textbf{{v}}}_{1}$. Let $V^{\prime}=V/R$ and choose a splitting $s\colon V^{\prime}\to V$ as in Section 9. Such a splitting is defined by a choice of $\delta\in V^{*}$ satisfying $\langle\delta,{\textbf{{v}}}_{1}\rangle=1$, then $s(V^{\prime})=\mathop{\mathrm{Ann}}\delta$. Let $\Gamma^{\prime}=\{\gamma^{\prime}_{1},\ldots,\gamma^{\prime}_{m}\}$ be the corresponding point configuration in $(V^{\prime})^{*}=\mathop{\mathrm{Ann}}R$, where $\gamma^{\prime}_{i}=s^{*}(\gamma_{i})=\gamma_{i}-\delta$.
@@
$({\textbf{{v}}}^{\prime},{\textbf{{z}}})$ $\mapsto{\textbf{{v}}}\cdot{\textbf{{z}}}=\bigl{[}e^{\langle\gamma^{\prime}_{1},{\textbf{{v}}}^{\prime}\rangle}z_{1}:\cdots:e^{\langle\gamma^{\prime}_{m},{\textbf{{v}}}^{\prime}\rangle}z_{m}\bigr{]}.$
Bosio [9] proves that this action is proper if and only if $(\mathcal{E},\Gamma^{\prime})$ satisfies the imbrication condition, such a proper action is free, and the quotient $P(\mathcal{E})/\widetilde{V}^{\prime}$ of a proper action is compact if and only if the substitute existence condition is satisfied.
@@
To link the definition of LVMB manifold with the holomorphic exponential actions considered in Section 10, we relate LVMB data $\{\mathcal{E},\Gamma^{\prime}\}$ to fan data. Let $\mathcal{K}$ be the pure $(m-k-1)$-dimensional simplicial complex whose maximal simplices are complements to $J\in\mathcal{E}$, that is,
@@
    $\{\mathcal{E},\Gamma^{\prime}\}$ is an LVMB datum if and only if $\{\mathcal{K},\mathrm{A}\}$ defines a complete simplicial fan;
@@
    the LVMB manifold $\mathcal{N}=P(\mathcal{E})/\widetilde{V}^{\prime}$ is the quotient of the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ by the diagonal free action of the circle $S^{1}$, and the complex structure is the one of the quotient of $U_{\mathbb{C}}(\mathcal{K})$ by the holomorphic exponential action of ××$\mathbb{C}^{\times}\times\widetilde{V}^{\prime}$.
@@
To prove (2), first observe that the imbrication condition is equivalent to the condition that $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$ and $\mathop{\mathrm{cone}}\Gamma_{\widehat{J}}$ have a common interior point for any $I,J\in\mathcal{K}$ (compare Theorem 9.2). The latter condition is equivalent to that $\{\mathcal{K},\mathrm{A}\}$ defines a simplicial fan by Proposition 4.4. Now the fan is complete if and only if the substitute existence condition is satisfied, see Proposition 6.2.
@@
Let $\{\mathcal{K},\mathrm{A}\}$ be the fan data corresponding to $\mathcal{N}$. We add a ghost vertex to $\mathcal{K}$. The data $\{\mathcal{K}_{+},\mathrm{A}_{+}\}$, where $\mathrm{A}_{+}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m},\bf 0\}$, defines the same fan, see Construction 10.8. We have $V_{+}=V\oplus\mathbb{R}\langle{\textbf{{v}}}_{+}\rangle=V^{\prime}\oplus R\oplus\mathbb{R}\langle{\textbf{{v}}}_{+}\rangle$. Then $Q=R\oplus\mathbb{R}\langle{\textbf{{v}}}_{+}\rangle$ is a two-dimensional rational subspace of $V_{+}$. It has a complex structure, which together with $\mathcal{J}^{\prime}\colon V^{\prime}\to V^{\prime}$ defines a complex structure on $V_{+}$. Denote the resulting complex spaces by $\widetilde{V}_{+}=\widetilde{V}^{\prime}\oplus\widetilde{Q}$. Then we have a diagram of holomorphic quotients (12.4):
@@
An LVM manifold is an LVMB manifold whose corresponding data $\{\mathcal{K},\mathrm{A}\}$ defines a normal fan. The original construction [27] is as follows.
@@
[weak hypebolicity] if $\mathbf{0}\in\mathop{\mathrm{conv}}\Gamma^{\prime}_{J}$, then $|J|\geqslant k$.
@@
It is shown in [27] that the exponential action $\widetilde{V}^{\prime}\times P(\mathcal{E})\to P(\mathcal{E})$ given by (13.1) is free and the quotient $P(\mathcal{E})/\widetilde{V}^{\prime}$ is identified with the nondegenerate intersection of projective quadrics
@@
Now let $V=V^{\prime}\oplus\mathbb{R}$, $\gamma_{i}=\begin{pmatrix}\gamma^{\prime}_{i}\\ 1\end{pmatrix}$, let $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ be the Gale dual configuration to $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$, and let $\mathcal{K}$ be the simplicial complex given by (13.2). By the holomorphic versions of Theorems 8.7 and 9.3, the data $\{\mathcal{K},\mathrm{A}\}$ defines the normal fan of a simple polytope $P$ given by (8.1), and and the quotient $U_{\mathbb{C}}(\mathcal{K})/(\mathbb{R}_{>}\times\widetilde{V}^{\prime})$ is diffeomorphic to the nondegenerate intersection of Hermitian quadrics
@@ 14. Toric varieties and irrational deformations
Toric varieties corresponding to complete simplicial rational fans can be obtained as partial quotients $\mathcal{Z}_{\mathcal{K}}/T_{V}$; this corresponds to the case when the entire space $V$ is rational with respect to the vector configuration $\Gamma$. By deforming the configuration $\Gamma$ we obtain holomorphic foliated manifolds $(\mathcal{Z}_{\mathcal{K}},\mathcal{F})$ which model irrational deformations of toric varieties.
@@ 14.1. Toric varieties as partial quotients
The Gale dual rational vector configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$ spans a full rank lattice $M^{*}$, which denote by $N$.
@@
The torus-exponential action (12.2) becomes the algebraic action of $\mathbb{C}^{\times}_{L}$ on $\mathbb{C}^{m}$. Given a simplicial complex $\mathcal{K}$ on $[m]$, we obtain an algebraic action of $\mathbb{C}^{\times}_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ by restriction. By Theorem 12.2, this action is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data, in which case the action is almost free with finite stabilizers.
The (geometric) quotient $U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ by a proper algebraic action of $\mathbb{C}^{\times}_{L}$ is the toric variety $X_{\Sigma}$ corresponding to the rational simplicial fan $\Sigma$ defined by $\{\mathcal{K},\mathrm{A}\}$ [15, §5.1]. The algebraic torus (×$(\mathbb{C}^{\times})^{m}/\mathbb{C}^{\times}_{L}=N\otimes_{\mathbb{Z}}\mathbb{C}^{\times}\cong(\mathbb{C}^{\times})^{n}$ acts on $X_{\Sigma}=U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ with a Zariski open orbit.
Remark.
The construction of toric varieties as quotients takes as input a lattice $N$ and a rational fan $\Sigma$ in the space $N\otimes_{\mathbb{Z}}\mathbb{R}$, which is $W^{*}$ in our notation. The vectors ${\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}$ are the primitive lattice vectors generating the one-dimensional cones of $\Sigma$. If the fan $\Sigma$ is simplicial, the toric variety $X_{\Sigma}$ can be defined as the geometric quotient $U_{\mathbb{C}}(\mathcal{K})/G$, where $G$ is the algebraic subroup of (×$(\mathbb{C}^{\times})^{m}$ given by (14.3). It may turn out that the vectors ${\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}$ do not span the lattice $N$; in particular, $\mathbb{Z}\langle{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\rangle$ may be a sublattice of finite index in $N$. It the latter case the group (14.3) is a product of an algebraic torus and a finite group $N/\mathbb{Z}\langle{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\rangle$. On the other hand, when we take as input a pair of Gale dual rational vector configurations $\Gamma$ and $\mathrm{A}$, the only natural choice for a lattice $N$ is $\mathbb{Z}\langle{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\rangle$, so the group (14.3) does not have finite factors and is isomorphic to the algebraic torus $\mathbb{C}^{\times}_{L}$.
@@
A complex structure on $\mathcal{Z}_{\mathcal{K}}$ is defined by a fan data $\{\mathcal{K},\mathrm{A}\}$ and a choice of complex structure on $V$, see §10. When the fan $\Sigma$ defined by $\{\mathcal{K},\mathrm{A}\}$ is rational, there is a holomorphic torus fibre bundle $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ over the corresponding toric variety. A perturbation of the vector configuration $\mathrm{A}$ destroys the rationality, the subgroup (14.2) ceases to be closed in (×$(\mathbb{C}^{\times})^{m}$, the closed holomorphic tori in the fibres of the bundle $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ “open up”, and the fibre bundle turns into a holomorphic foliation $\mathcal{F}$ of $\mathcal{Z}_{\mathcal{K}}$ with noncompact leaves $G/\widetilde{V}$. The holomorphic foliated manifolds $(\mathcal{Z}_{\mathcal{K}},\mathcal{F})$ therefore model irrational deformations of toric varieties, which have recently been studied from several different perspectives in the works [5, 7, 33, 24, 23, 6].

### v2

@@ 13. LVM and LVMB manifolds
We start with a spanning vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$ for which there exists ${\textbf{{v}}}_{1}\in V$ such that $\langle\gamma_{i},{\textbf{{v}}}_{1}\rangle=1$ for $i=1,\ldots,m$. Equivalently, the Gale dual configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ satisfies the relation $\sum_{k=1}^{m}{\textbf{{a}}}_{i}=\bf 0$. In particular, the conditions of Proposition 9.1 are met. Let $R=\mathbb{R}\langle{\textbf{{v}}}_{1}\rangle$. Thinking in terms of relations between the vectors of $\mathrm{A}$, the subspace $R$ is generated by a single relation with all coefficients $1$, so it is rational. The lattice (11.1) is the $\mathbb{Z}$-span of ${\textbf{{v}}}_{1}$. Let $V^{\prime}=V/R$ and choose a splitting $s\colon V^{\prime}\to V$ as in Section 9. Such a splitting is defined by a choice of $\delta\in V^{*}$ satisfying $\langle\delta,{\textbf{{v}}}_{1}\rangle=1$, then $s(V^{\prime})=\mathop{\mathrm{Ann}}\delta$. Let $\Gamma^{\prime}=\{\gamma^{\prime}_{1},\ldots,\gamma^{\prime}_{m}\}$ be the corresponding point configuration in $(V^{\prime})^{*}=\mathop{\mathrm{Ann}}R$, where $\gamma^{\prime}_{i}=s^{*}(\gamma_{i})=\gamma_{i}-\delta$.
@@
$({\textbf{{v}}}^{\prime},{\textbf{{z}}})$ $\longmapsto{\textbf{{v}}}\cdot{\textbf{{z}}}=\bigl{[}e^{\langle\gamma^{\prime}_{1},{\textbf{{v}}}^{\prime}\rangle_{\mathbb{C}}}z_{1}:\cdots:e^{\langle\gamma^{\prime}_{m},{\textbf{{v}}}^{\prime}\rangle_{\mathbb{C}}}z_{m}\bigr{]}.$
Bosio proves in [10, Theorem 1.4] that this action is proper if and only if $(\mathcal{E},\Gamma^{\prime})$ satisfies the imbrication condition, such a proper action is free, and the quotient $P(\mathcal{E})/\widetilde{V}^{\prime}$ of a proper action is compact if and only if the substitute existence condition is satisfied.
@@
To link the definition of LVMB manifold with the holomorphic exponential actions considered in Section 10, we relate LVMB data $\{\mathcal{E},\Gamma^{\prime}\}$ to fan data (triangulated vector configurations). Let $\mathcal{K}$ be the pure $(m-k-1)$-dimensional simplicial complex whose maximal simplices are complements to $J\in\mathcal{E}$, that is,
@@
    $\{\mathcal{E},\Gamma^{\prime}\}$ is an LVMB datum if and only if $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration defining a complete simplicial fan;
@@
    the LVMB manifold $\mathcal{N}=P(\mathcal{E})/\widetilde{V}^{\prime}$ is the quotient of the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ by the diagonal free action of the circle $S^{1}$, and the complex structure is the one of the quotient of $U_{\mathbb{C}}(\mathcal{K})$ by the holomorphic torus-exponential action of ××$\mathbb{C}^{\times}\times\widetilde{V}^{\prime}$.
@@
To prove (2), first observe that the imbrication condition is equivalent to the condition that $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$ and $\mathop{\mathrm{cone}}\Gamma_{\widehat{J}}$ have a common interior point for any $I,J\in\mathcal{K}$ (compare Theorem 9.2). The latter condition is equivalent to $\{\mathcal{K},\mathrm{A}\}$ being a triangulated configuration by Proposition 4.4. Now the fan is complete if and only if the substitute existence condition is satisfied, see Proposition 6.2.
@@
Let $\{\mathcal{K},\mathrm{A}\}$ be the triangulated configuration corresponding to $\mathcal{N}$. We add a ghost vertex to $\mathcal{K}$. The triangulated configuration $\{\mathcal{K}_{+},\mathrm{A}_{+}\}$, where $\mathrm{A}_{+}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m},\bf 0\}$, defines the same fan, see Construction 10.8. We have $V_{+}=V\oplus\mathbb{R}\langle{\textbf{{v}}}_{+}\rangle=V^{\prime}\oplus R\oplus\mathbb{R}\langle{\textbf{{v}}}_{+}\rangle$. Then $Q=R\oplus\mathbb{R}\langle{\textbf{{v}}}_{+}\rangle$ is a two-dimensional rational subspace of $V_{+}$. It has a complex structure, which together with $\mathcal{J}^{\prime}\colon V^{\prime}\to V^{\prime}$ defines a complex structure on $V_{+}$. Denote the resulting complex spaces by $\widetilde{V}_{+}=\widetilde{V}^{\prime}\oplus\widetilde{Q}$. Then we have a diagram of holomorphic quotients (12.4):
@@
An LVM manifold is an LVMB manifold whose corresponding data $\{\mathcal{K},\mathrm{A}\}$ define a normal fan. The original construction [34] is as follows.
@@
[weak hyperbolicity] if $\mathbf{0}\in\mathop{\mathrm{conv}}\Gamma^{\prime}_{J}$, then $|J|\geqslant k$.
@@
It is shown in [34] that the holomorphic exponential action $\widetilde{V}^{\prime}\times P(\mathcal{E})\to P(\mathcal{E})$ given by (13.1) is free and the quotient $P(\mathcal{E})/\widetilde{V}^{\prime}$ is identified with the nondegenerate intersection of projective quadrics
@@
Now let $V=V^{\prime}\oplus\mathbb{R}$, $\gamma_{i}=\begin{pmatrix}\gamma^{\prime}_{i}\\ 1\end{pmatrix}$, let $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ be the Gale dual configuration to $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$, and let $\mathcal{K}$ be the simplicial complex given by (13.2). By the holomorphic versions of Theorems 8.7 and 9.3, the data $\{\mathcal{K},\mathrm{A}\}$ define the normal fan of a simple polytope $P$ given by (8.1), and and the quotient $U_{\mathbb{C}}(\mathcal{K})/(\mathbb{R}_{>}\times\widetilde{V}^{\prime})$ is diffeomorphic to the nondegenerate intersection of Hermitian quadrics
@@
LVM manifolds can also be characterised within LVMB manifolds in terms of the canonical foliation $\mathcal{F}$ on an LVMB manifold, introduced by Cupit-Foutou and Zaffran in [17]. The foliation $\mathcal{F}$ turns into a holomorphic fibration over a toric orbifold with fibres compact tori in the rational case. Ishida [24] proves that the canonical foliation $\mathcal{F}$ is transversely Kähler if and only if the LVMB manifold is LVM.
Complex geometry of moment-angle manifolds and their partial quotients, including LVM- and LVMB-manifolds, is very rich and far from being completely understood. Analytic subsets and invariants of the complex structure of these manifolds, including the Hodge numbers and Dolbeault cohomology, holomorphic automorphism groups and moduli spaces of complex structures are currently the subject of active research. We refer to [35, 11, 38, 43, 9, 15, 39, 30, 28] for the aspects of complex geometry of these manifolds.
@@ 14. Toric varieties and irrational deformations
Complete nonsingular toric varieties are defined by nonsingular rational fans and can be obtained as partial quotients $\mathcal{Z}_{\mathcal{K}}/T_{V}$; this corresponds to the case when the entire space $V$ is rational with respect to the vector configuration $\Gamma$. By deforming the configuration $\Gamma$ we obtain holomorphic foliated manifolds $(\mathcal{Z}_{\mathcal{K}},\mathcal{F})$ which model irrational deformations of simplicial toric varieties.
@@ 14.1. Toric varieties as partial quotients
The Gale dual rational vector configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$ spans a full rank lattice $M^{*}$, which we denote by $N$.
@@
The torus-exponential action (12.2) becomes the algebraic action of $\mathbb{C}^{\times}_{L}$ on $\mathbb{C}^{m}$. Given a simplicial complex $\mathcal{K}$ on $[m]$, we obtain an algebraic action of $\mathbb{C}^{\times}_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ by restriction. By Theorem 12.2, this action is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, in which case the action is almost free with finite stabilisers.
When the vectors of the configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ are primitive in the lattice $N$ spanned by $\mathrm{A}$, the (geometric) quotient $U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ by a proper algebraic action of $\mathbb{C}^{\times}_{L}$ is the toric variety $X_{\Sigma}$ corresponding to the rational simplicial fan $\Sigma$ defined by $\{\mathcal{K},\mathrm{A}\}$ [16, §5.1]. The algebraic torus (×$(\mathbb{C}^{\times})^{m}/\mathbb{C}^{\times}_{L}=N\otimes_{\mathbb{Z}}\mathbb{C}^{\times}\cong(\mathbb{C}^{\times})^{n}$ acts on $X_{\Sigma}=U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ with a Zariski open orbit. The toric variety $X_{\Sigma}$ is nonsingular if $\{\mathcal{K},\mathrm{A}\}$ is a nonsingular triangulated configuration (this implies that the action of $\mathbb{C}^{\times}_{L}$ is free and the vectors ${\textbf{{a}}}_{i}$ are primitive in the lattice $N$).
More generally, when the vectors ${\textbf{{a}}}_{i}$ are not necessarily primitive in the lattice $N$, the quotient $U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ is still the toric variety $X_{\Sigma}$ corresponding to the fan $\Sigma$ defined by $\{\mathcal{K},\mathrm{A}\}$, yet with the non-standard orbifold structure. This phenomenon is treated in [35, §4], see also [8]. We illustrate it on a simple example.
Example 14.1.
Consider the two-vector configuration $\mathrm{A}=\{2,-1\}$ in one-dimensional space $W^{*}=\mathbb{R}$ and $\mathcal{K}=\{\varnothing,\{1\},\{2\}\}$. Then the triangulated configuration $\{\mathcal{K},\mathrm{A}\}$ defines a complete nonsingular fan in $\mathbb{R}$, but the vector ${\textbf{{a}}}_{1}=2$ is non-primitive in the lattice $N=\mathbb{Z}$ spanned by $\mathrm{A}$. The Gale dual configuration is $\Gamma=\{1,2\}$ and the group (14.2) is given by
$G=\{(e^{v},e^{2v})=(t,t^{2})\in(\mathbb{C}^{\times})^{2}\}.$
The $G$-action on $U_{\mathbb{C}}(\mathcal{K})=\mathbb{C}^{2}\setminus\{0\}$ has a nontrivial stabiliser of order $2$ at the point $(0,1)$. The quotient $\mathbb{C}^{2}\setminus\{0\}/G$ is the projective line $\mathbb{C}P^{1}$ as a variety, but but comes equipped with an orbifold structure with one singular point of index $2$.
@@
A complex structure on $\mathcal{Z}_{\mathcal{K}}$ is defined by a triangulated configuration $\{\mathcal{K},\mathrm{A}\}$ and a choice of complex structure on $V$, see §10. When the fan $\Sigma$ defined by $\{\mathcal{K},\mathrm{A}\}$ is rational, there is a holomorphic torus fibre bundle $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ over the corresponding toric variety. A perturbation of the vector configuration $\mathrm{A}$ destroys the rationality, the subgroup (14.2) ceases to be closed in (×$(\mathbb{C}^{\times})^{m}$, the closed holomorphic tori in the fibres of the bundle $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ “open up”, and the fibre bundle turns into a holomorphic foliation $\mathcal{F}$ of $\mathcal{Z}_{\mathcal{K}}$ with noncompact leaves $G/\widetilde{V}$. Then $\mathcal{F}$ is an example of the canonical foliation on a complex manifold, considered in [25] (see also [28]). The holomorphic foliated manifolds $(\mathcal{Z}_{\mathcal{K}},\mathcal{F})$ therefore model irrational deformations of toric varieties, which have recently been studied from several different perspectives in the works [5, 7, 41, 29, 28, 6].
On the other hand, the construction of toric varieties as quotients takes as input a lattice $N$ and a rational fan $\Sigma$ in the space $N\otimes_{\mathbb{Z}}\mathbb{R}$, which is $W^{*}$ in our notation. The vectors ${\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}$ are the primitive lattice vectors generating the one-dimensional cones of $\Sigma$. If the fan $\Sigma$ is simplicial, the toric variety $X_{\Sigma}$ can be defined as the geometric quotient $U_{\mathbb{C}}(\mathcal{K})/G$, where $G$ is the algebraic subgroup of (×$(\mathbb{C}^{\times})^{m}$ given by (14.3). It may turn out that the vectors ${\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}$ do not span the lattice $N$; in particular, $\mathbb{Z}\langle{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\rangle$ may be a sublattice of finite index in $N$. In the latter case the group (14.3) is a product of an algebraic torus and a finite group $N/\mathbb{Z}\langle{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\rangle$. When we take as input a pair of Gale dual rational vector configurations $\Gamma$ and $\mathrm{A}$, the only natural choice for a lattice $N$ is $\mathbb{Z}\langle{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\rangle$, so the group (14.3) does not have finite factors and is isomorphic to the algebraic torus $\mathbb{C}^{\times}_{L}$. To take account of all simplicial toric varieties, one can follow the approach of Prato [40] to toric quasifolds and include a finitely generated abelian subgroup $N$ of $W^{*}$ (a quasilattice) containing the vectors ${\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}$ as part of the input data. In this way, all complete simplicial toric varieties (orbifolds), as well as toric quasifolds, can be obtained as quotients of LVMB manifolds [8, Theorem 3.4].